On July 1, 2019, Sugarland Company issued $2,000,000 face value of 10%, 10-year bonds at $1,770,602, a price which implies an effective interest rate of 12%. Sugarland uses the effective-interest method to amortize bond premiums and discounts. These bonds pay interest semiannually on June 30 and December 31.

Required: Compute the answers to the following questions:

(a.) How much interest will Sugarland pay (in cash) every six months?

(b.)What is the dollar amount of the premium or discount on these bonds at the date of issue? (Indicate whether it is a premium or discount.)

(c.) How much interest expense will Sugarland recognize for the six months ended December 31, 2019?

(d.)How much interest expense will Sugarland recognize for the six months ended June 30, 2020?
Answer #1

a) Interest paid every six month = 2000000*10%*6/12 = 100000

b) Discount on bonds payable = 2000000-1770602 = 229398

c) Interest expense on december 31,2019 = 1770602*6% = 106236.12 or 106236

d) Interest expense on June 30,2020 = (1770602+6236)*6% = 106610